### v1

@@ 4.1 Proposed testing procedure
As demonstrated in Section 3 and Appendix A, strong collusive firm behavior, if properly accounted for, increases the power of the own-firm IVs, $\mathbf{z}^{own}$, in the first-stage price regression. More specifically, the product attributes of another firm unexpectedly serve as stronger instruments for a firm’s own product prices if that firm is indeed a colluding partner. In such cases, including these attributes in the own-firm instruments may enhance the strength of the instruments. Conversely, their inclusion would weaken the instruments’ strength in the first-stage price regression if the two firms are actually engaging in price competition.
Based on this intuition, we propose that collusive behaviors among firms can be assessed by comparing the IV performance in the two first-stage price regressions. Specifically, we construct two distinct sets of BLP-style instruments: one based on the observed firm index (competition IVs), denoted by $\mathbf{z}^{comp}$, and the other based on the suspected colluding firm index (collusion IVs), denoted by $\mathbf{z}^{coll}$. We then estimate two separate first-stage regressions of prices using these two sets of instruments individually, and compare their relative strengths using the non-nested model selection method proposed by rivers2002model. [^8] The implementation of rivers2002model’s test in cross-sectional and panel regression settings can be inferred from wooldridge2010 (Section 13.11.2).
@@
Note that $\mathscr{J}_{f_{c}t}$ is a set of products owned by firm $f$ and its suspected colluding partners, as $f_{c}$ indexes firm $f$ and its suspected colluding partners together. The choice and functional form of instruments in (8) are selected for brevity and clarity, with further discussion provided in Section 6.3 and Appendix B.
We emphasize that $\mathbf{z}_{jt}^{comp}$ is constructed based on the observed competitive market ownership, whereas $\mathbf{z}_{jt}^{coll}$ is constructed based on the suspected competitive market ownership. For example, suppose that the regulatory authority suspects that, among the four firms in a product market, indexed as A, B, C, and D, two firms, A and B, are engaging in price collusion. Based on observed price and product attributes data, the authority can construct two sets of BLP-type instruments as described above. In this case, $\mathbf{z}_{jt}^{comp}$ treats firms A and B as two separate entities, whereas $\mathbf{z}_{jt}^{coll}$ treat these two firms as a single entity, reflecting their suspected collusion.
Next, we run the following two separate first-stage price regressions using $\mathbf{z}_{jt}^{comp}$ and $\mathbf{z}_{jt}^{coll}$:
@@
where $x_{jt}$ represents the exogenous characteristics of product $j$, including an intercept. Continuing with the previous example, if firms A and B indeed collude and internalize the profit of their colluding partner, then $\mathbf{z}_{jt}^{coll}$ should explain the equilibrium prices better than $\mathbf{z}^{comp}$ do. On the other hand, if the two firms do not engage in collusion and all four firms compete à la Bertrand, then $\mathbf{z}^{comp}$ should explain the equilibrium prices better.
@@ 4.2 Existing Testing Procedure
where $x_{jt}$ and $w_{jt}$ represent observed product attributes and cost shifters excluded from the demand function, respectively, while $\omega_{jt}$ denotes an unobserved cost component. To convey the main idea and simplify discussion, we impose the following two assumptions: (i) the marginal cost function, $h(x_{jt},w_{jt})$, which must be pre-specified by researchers to implement the testing procedure, is linear in $x_{jt}$ and $w_{jt}$, and (ii) marginal cost is constant with respect to quantities produced, implying no economies of scale or scope.999backus2021common proposed a testing procedure that incorporates a flexible specification of the cost function. Additionally, the marginal cost can be nonconstant with respect to quantities produced. In such cases, the instruments are required to satisfy additional conditions to distinguish between alternative firm conduct models (berry2014identification; backus2021common; duarte2024testing; dearing2024learning).
@@
where $\omega_{jt}^{0}=p_{jt}-\eta_{jt}^{0}-h(x_{jt},w_{jt})$. A vector of excluded instruments, denoted by $\mathbf{z}_{jt}$, typically consists of BLP-style instruments (functions of exogenous product attributes) and cost shifters excluded from the demand specification, $w_{jt}$.101010backus2021common proposed a functional form for BLP-style instruments to enhance the power of the RV testing framework, drawing on the literature on optimal instruments in nonlinear GMM settings (chamberlain1987asymptotic). This functional form results in a scalar moment that is independent of the choice of the weighting matrix used to form the objective function. Additionally, the choice of instruments (beyond BLP-style instruments) to satisfy falsifiable conditions to distinguish between two alternative conduct models affects the validity of testing procedure. See dearing2024learning for discussions on falsifiable conditions and empirical applications.
@@ 5.1 Data-generating process
There are two exogenous product characteristics, $1$ and $x_{jt}$, and one exogenous cost shifter, $w_{jt}$, where both $x_{jt}$ and $w_{jt}$ are randomly drawn from a standard uniform distribution. Price ($p_{jt}$) is an endogenous product characteristic determined by equilibrium conditions in a differentiated goods market under a Bertrand-Nash framework. The true linear demand parameters are given by $\alpha=1$, $\beta_{1}=-4.5$, and $\beta_{2}=6$. The heterogeneous component of demand, when included, is given by $\sigma_{x}v_{i}x_{jt}$, where $\sigma_{x}=3$ and $v_{i}$ is the standard normal random variable. Cost parameters are given by $\gamma_{1}=2$, $\gamma_{2}=1$, and $\gamma_{3}=0.2$. The unobservable components of demand and cost, $(\xi_{jt},\omega_{jt})$, are randomly drawn from a mean-zero bivariate normal distribution with standard deviations $\sigma_{\xi}=0.2$, $\sigma_{\omega}=0.2$, and covariance $\sigma_{\xi\omega}=0.1$.
@@
Firms may engage in collusion, internalizing the profits of their colluding partners as part of their own. The degree of collusion is measured by $F/F_{c}$, where $F_{c}\in\{1,\ldots,F-2,F-1\}$ denotes the effective number of competitors – i.e., the number of distinct competitive entities after treating all colluding firms as a single entity. We assume that the first $F-F_{c}+1$ firms collude or are suspected of colluding by regulatory authorities. For example, suppose $F=6$ and $F_{c}=4$. In this case, the degree of collusion is 1.5, with the first three firms colluding to jointly set profit-maximizing prices, leaving the effective competitors indexed as $f_{c}=1,4,5,6$. Regulatory authorities suspect three firms ($f=1,2,3$) of engaging in collusion. The degree of internalization, represented by the conduct parameter $\phi\in[0,1]$, quantifies the weight assigned to the profits of colluding partners, as shown in equation (3). Continuing with the previous example, when $F_{c}=4$ but $\phi=0$, this configuration represents a scenario in which regulatory authorities falsely suspect three firms of collusion when they are actually engaged in price competition.
@@
Equilibrium prices in market $t$, $p_{t}$, are endogenously determined by the Bertrand first-order conditions (4). These prices are solved using the fixed-point algorithm proposed by morrow2011fixed and implemented in PyBLP (conlon2020best). We use 9 Gauss-Hermite quadrature nodes to numerically integrate choice probabilities when a random coefficient $\sigma_{x}v_{i}$ is included in the indirect utility function.
@@
We also consider the case where the product attribute is correlated with the unobservable demand component $\xi_{jt}$, and is therefore endogenous. We denote the endogenous attribute as $x_{jt}^{endo}$ and derive it as follows. First, we compute $x_{jt}^{unscaled}=x_{jt}+\rho\cdot\xi_{jt}$, where $\rho\in\{-10,-5,-1,0,1,5,10\}$ determines the degree and direction of endogeneity; the case $\rho=0$ corresponds to an exogenous product attribute. Next, given that the variation of $x_{jt}^{unscaled}$ depends on $\rho$, we apply min-max normalization to define $x_{jt}^{endo}=\frac{x_{jt}^{unscaled}-x_{t}^{\min}}{x_{t}^{\max}-x_{t}^{\min}}$, where $x_{t}^{\min}$ and $x_{t}^{\max}$ are the minimum and maximum values of $x_{jt}^{unscaled}$ in market $t$. In this way, we fix the support of the endogenous attribute: $x_{jt}^{endo}\in[0,1]$ for any $\rho$.
@@
Each Monte Carlo configuration is represented by a unique five-tuple $(J,F,T,\phi,F_{c})$. Specifically, $J=36$, $F\in\{1,2,\ldots,36\}$, $T\in\{10,100\}$, $\phi\in\{0,0.1,\ldots,1\}$, and $F_{c}\in\{1,2,\ldots,F-1\}$. [^14] When the DGP includes an endogenous product attribute, $\phi$ is fixed at 1, and $\rho$ is varied to generate unique five-tuples $(J,F,T,\rho,F_{c})$. For each configuration, we construct 500 simulated datasets ($S=500$) to evaluate the finite-sample performance of the proposed test statistic ($T_{IV}^{RV}$) and compare it with the performance of the test based on equilibrium markup conditions ($T_{markup}^{RV}$). We also investigate the performance of $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as instruments for demand estimation. Furthermore, in Section 7, we focus on cases with varying $F$ and fixed $J$ but without collusive behavior ($\phi=0$ and $F_{c}=F$) to evaluate the performance of BLP-type IVs, following the approach of armstrong2016large.
@@ 5.2 Computation of test statistics
For each dataset with configuration $(J,F,T,\phi,F_{c})$, we construct $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as outlined in Section 4.1.151515Note that the summation of the constant term, often referred to as the product counts IV, cannot be utilized in our DGP setting because it fails to generate cross-firm or cross-market variations in the number of products. Moreover, the inclusion of second-order polynomial IVs in (8) ensures that the rank condition for the GMM framework is satisfied in our DGP when a random coefficient is incorporated into the indirect utility function (15). Note that even when $\phi=0$, if $F_{c}<F$, we still construct $\mathbf{z}^{coll}$ based on the scenario in which the researcher has a false suspicion of collusion among the firms of interest. To evaluate the relevance of these two instruments under a given true conduct model, we calculate two F-statistics associated with (i) $\bm{\theta}^{1}=0$ (Fstat1) and (ii) $\bm{\theta}^{2}=0$ (Fstat2) from equations (9) and (10), respectively. Our proposed test statistic $T_{IV}^{RV}$ is then computed using equation (11) and the description provided therein. Heteroskedasticity-robust standard errors are used to compute $\hat{\sigma}_{IV}$ in equation (11).
@@
as excluded instruments $\mathbf{z}_{jt}=\left(\mathbf{z}_{jt}^{own},\mathbf{z}_{jt}^{other}\right)$. Note that these instruments are constructed based on observed ownership rather than suspected ownership. Next, we specify the marginal cost function as a linear function of the constant term and $x_{jt}$ only, even though it is also linear in a cost shifter $w_{jt}$, as shown in equation (16). [^16] Accordingly, we do not include $w_{jt}$ as an instrument when estimating the demand. Thus, we consider a scenario in which the cost function is potentially misspecified, either due to data limitations on cost shifters or the omission of a relevant cost shifter. This approach enables us to evaluate the performance of the two testing procedures without using additional data. We then follow the procedure outlined in Section 4.2 to construct $T_{markup}^{RV}$ under two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full internalization under the ownership structure consistent with the true effective competitors index used in the DGP). [^17] Although the testing procedure built on equilibrium markup conditions allows us to compare conduct models with other values of $\phi$, such as $\phi=0.2$ versus $\phi=0.8$, the primary interest of researchers or regulatory authorities is often to test full collusion ($\phi=1$) against competition ($\phi=0$). Therefore, we compute $T_{markup}^{RV}$ under these two conduct models. The standard error $\hat{\sigma}_{markup}$ in equation (14) is adjusted for two-step demand estimation errors and is heteroskedasticity-robust. The computation of $T_{markup}^{RV}$ is done using PyRVtest (pyrvtest).
6 Monte Carlo evidence: testing performance of $T_{IV}^{RV}$ vs. $T_{markup}^{RV}$
@@ 6.1 Power of instruments and estimation performance
We begin by preliminarily assessing the power of our proposed testing procedure outlined in Section 4.1. To do so, we consider a scenario with four firms, two of which collude ($F=4,F_{c}=3$) with the degree of internalization of their colluding partners’ profits denoted by $\phi$. We generate Monte Carlo datasets, as outlined in Section 5, for various values of $\phi\in[0,1]$. The top panel of Table 1 presents the results when a random coefficient is excluded from the indirect utility function (15) in the DGP, while the bottom panel shows the results when it is included in the utility function.
@@
Table 1: Comparison of IV performance: $\mathbf{z}^{comp}$ vs. $\mathbf{z}^{coll}$ ($F=4$, $F_{c}=3$, $T=100$)
@@
Internalization Median values ($S=500$) of % of Fstat2 Median
degree ($\phi$) $|\alpha-\hat{\alpha}|$ $|\sigma_{x}-\hat{\sigma}_{x}|$ $\text{Fstat}_{1}$ $|\alpha-\hat{\alpha}|$ $|\sigma_{x}-\hat{\sigma}_{x}|$ $\text{Fstat}_{2}$ $>$ Fstat1 of $s_{o}$
@@
Notes: The table compares the median values of the absolute error of the estimated price coefficient, the absolute error of the estimated nonlinear coefficient (reported in the bottom panel), and the F-statistics across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=4,T=100,\phi,F_{c}=3$), using $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as instruments individually. The top panel presents the results when a random coefficient is excluded from the indirect utility function (15) in the DGP, while the bottom panel presents the results when it is included in the utility function.
Our simulation results presented in Table 1 also show that the median absolute error of the estimated price coefficient $\hat{\alpha}$ decreases significantly in response to a rise in the degree of internalization when we use $\mathbf{z}^{coll}$ as instruments. As a result, $\mathbf{z}^{coll}$ outperforms $\mathbf{z}^{comp}$ in estimating $\alpha$ when firms collude, even with a low degree of internalization ($\phi\geq 0.4$ without a random coefficient and $\phi\geq 0.1$ with a random coefficient).
A similar pattern is observed for the median absolute error of the estimated non-linear coefficient $\hat{\sigma}_{x}$. Moreover, for any $\phi$, the median absolute error obtained using $\mathbf{z}^{coll}$ as instruments is smaller than that obtained using $\mathbf{z}^{comp}$. This finding is not contradictory to our earlier observation that, when the degree of internalization is low ($\phi\leq 0.3$), using $\mathbf{z}^{coll}$ instead of $\mathbf{z}^{comp}$ leads to lower first-stage F-statistics. This is because identifying $\sigma_{x}$ requires instruments that are correlated not only with price but also with market shares. [^18] As a simple illustration, consider a nested logit specification, a type of random coefficient model. After inverting a demand system (Berry94), the resulting estimation equation using a standard linear instrumental variables approach is as follows: $\log\frac{s_{jt}}{s_{ot}}=\beta x_{jt}-\alpha p_{jt}+\sigma\log s_{j|g,t}+\xi_{jt},$ where $s_{j|g,t}$ denotes the market share of product $j$ within group $g$. Identifying this random coefficient model requires instruments for both the endogenous price and the endogenous market share $s_{j|g,t}$. See berry2014identification for the instrument requirements under general nonparametric settings. We revisit this point in Section 7, where we discuss the role of own- and other-firm instruments.

### v2

@@ 4.1 Proposed testing procedure
As demonstrated in Section 3 and Appendix LABEL:app:_partial_derivative, properly accounting for strong collusive firm behavior enhances the power of BLP-style instruments in the first-stage price regression. More specifically, when firms collude, incorporating the attributes of colluding partners into own-firm instruments strengthens the instruments. Conversely, if firms actually engage in price competition, this incorporation would weaken the instruments’ strength in the first-stage price regression.
Based on this intuition, we propose assessing collusive behaviors among firms by comparing the performance of IVs in two first-stage price regressions. Specifically, we construct two distinct sets of BLP-style instruments: one based on the observed firm index (competition IVs), denoted by $\mathbf{z}^{comp}$, and the other based on the suspected colluding firm index (collusion IVs), denoted by $\mathbf{z}^{coll}$. We then estimate two separate first-stage price regressions using these two sets of instruments individually and compare their relative strengths using the non-nested model selection method proposed by rivers2002model. [^12] The implementation of rivers2002model’s test in cross-sectional and panel regression settings can be inferred from wooldridge2010 (Section 13.11.2).
@@
Note that $\mathscr{J}_{f_{c}t}$ denotes the set of products owned by firm $f$ and its suspected colluding partners, as $f_{c}$ indexes these firms as a single entity. The choice of instruments and functional form in (9) is selected for brevity and clarity, with further discussion provided in Section 6.3 and Appendix LABEL:app:_functional_forms.
We emphasize that $\mathbf{z}_{jt}^{comp}$ is constructed based on the observed competitive market ownership, whereas $\mathbf{z}_{jt}^{coll}$ is constructed based on the suspected collusive market ownership. For example, suppose that the regulatory authority suspects that, among the four firms in a product market, indexed as $A$, $B$, $C$, and $D$, two firms, $A$ and $B$, are engaging in price collusion. Based on observed price and product attribute data, the authority can construct two sets of BLP-type instruments as described above. In this case, $\mathbf{z}_{jt}^{comp}$ treats firms $A$ and $B$ as two separate entities, whereas $\mathbf{z}_{jt}^{coll}$ treats these two firms as a single entity, reflecting their suspected collusion.
Next, we run the following two first-stage price regressions using $\mathbf{z}_{jt}^{comp}$ and $\mathbf{z}_{jt}^{coll}$ separately:
@@
where $x_{jt}$ represents the exogenous characteristics of product $j$, including an intercept. Continuing with the previous example, if firms $A$ and $B$ indeed collude and internalize the profit of their colluding partner, then $\mathbf{z}_{jt}^{coll}$ should explain the equilibrium prices better than $\mathbf{z}^{comp}_{jt}$ does. On the other hand, if the two firms do not engage in collusion and all four firms compete à la Bertrand, then $\mathbf{z}_{jt}^{comp}$ should explain the equilibrium prices better.
@@
It is important to note that our test allows for potential misspecification of competing models. More specifically, it does not require either model (10) or (11) to be the true pricing equation, nor do they need to represent structural pricing functions. [^13] In general, a structural pricing function cannot be estimated, as it emerges from the equilibrium behavior of firms and thus depends on numerous observed and unobserved factors. Instead, the RV test evaluates the relative fit of two alternative models, making our test both easy to implement and highly practical. This flexibility is a key advantage over other existing methods, which may require consistent estimates of the conditional mean function or structural components.
@@ 4.2 Existing Testing Procedure
where $x_{jt}$ and $w_{jt}$ represent observed product attributes and cost shifters excluded from the demand function, respectively, while $\omega_{jt}$ denotes an unobserved cost component. To simplify the discussion while conveying the main idea, we impose the following two assumptions: (i) the marginal cost function, $h(x_{jt},w_{jt})$, which must be pre-specified by researchers to implement the testing procedure, is linear in $x_{jt}$ and $w_{jt}$; and (ii) marginal cost is constant with respect to the quantity produced, implying no economies of scale or scope.141414backus2021common proposed a testing procedure that incorporates a flexible specification of the cost function. Additionally, the marginal cost can be nonconstant with respect to quantities produced. In such cases, the instruments are required to satisfy additional conditions to distinguish between alternative firm conduct models. See duarte2025conduct for details and applications.
@@
where $\omega_{jt}^{0}=p_{jt}-\eta_{jt}^{0}-h(x_{jt},w_{jt})$. A vector of excluded instruments, denoted by $\mathbf{z}_{jt}$, typically consists of BLP-style instruments (functions of exogenous product attributes) and cost shifters excluded from the demand specification, $w_{jt}$.151515backus2021common proposed a functional form for BLP-style instruments to enhance the power of the RV testing framework, drawing on the literature on optimal instruments in nonlinear GMM settings (chamberlain1987asymptotic). This functional form results in a scalar moment that is independent of the choice of the weighting matrix used to form the objective function. Additionally, the choice of instruments (beyond BLP-style instruments) to satisfy falsifiable conditions to distinguish between two alternative conduct models affects the validity of testing procedure. See dearing2024learning for discussions on falsifiable conditions and empirical applications.
@@ 5.1 Data-generating process
There are two exogenous product characteristics – a constant term and $x_{jt}$ – and one exogenous cost shifter, $w_{jt}$, where both $x_{jt}$ and $w_{jt}$ are randomly drawn from a standard uniform distribution. Price, $p_{jt}$, is an endogenous product characteristic determined by equilibrium conditions in a differentiated goods market under a Bertrand-Nash framework. The true linear demand parameters are given by $\alpha=1$, $\beta_{1}=-4.5$, and $\beta_{2}=6$. When included, the heterogeneous component of demand is given by $\sigma_{x}v_{i}x_{jt}$, where $\sigma_{x}=3$ and $v_{i}$ follows a standard normal distribution. Cost parameters are given by $\gamma_{1}=2$, $\gamma_{2}=1$, and $\gamma_{3}=0.2$. The unobservable components of demand and cost, $(\xi_{jt},\omega_{jt})$, are randomly drawn from a mean-zero bivariate normal distribution with standard deviations $\sigma_{\xi}=0.2$, $\sigma_{\omega}=0.2$, and covariance $\sigma_{\xi\omega}=0.1$.
@@
Firms may engage in collusion, internalizing the profits of their colluding partners as part of their own. The degree of collusion is measured by $F/F_{c}$, where $F_{c}\in\{1,\ldots,F-2,F-1\}$ denotes the effective number of competitors – i.e., the number of distinct competitive entities after treating all colluding firms as a single entity. We assume that the first $F-F_{c}+1$ firms either collude or are suspected of colluding by the regulatory authority. For example, suppose $F=6$ and $F_{c}=4$. In this case, the degree of collusion is 1.5, with the first three firms ($f=1,2,3$) colluding or suspected of colluding to jointly set profit-maximizing prices, leaving the effective competitors indexed as $f_{c}=1,4,5,6$. The degree of profit internalization, represented by the conduct parameter $\phi\in[0,1]$, quantifies the weight assigned to the profits of colluding partners, as shown in equation (4). Continuing with the previous example, when $F_{c}=4$ but $\phi=0$, this configuration represents a scenario in which regulatory authorities falsely suspect three firms of collusion when, in fact, they are engaged in price competition.
@@
Equilibrium prices in each market are endogenously determined by the Bertrand first-order conditions (5). These prices are solved using the fixed-point algorithm proposed by morrow2011fixed and implemented in PyBLP (conlon2020best). We use nine Gauss-Hermite quadrature nodes to numerically integrate the choice probabilities when a random coefficient is included in the indirect utility function.
@@
We also consider the case in which the product attribute is correlated with the unobservable demand component, $\xi_{jt}$, making it endogenous. We denote the endogenous attribute as $x_{jt}^{endo}$ and derive it as follows. First, we compute $x_{jt}^{unscaled}=x_{jt}+\rho\cdot\xi_{jt}$, where $\rho\in\{-10,-5,-1,0,1,5,10\}$ determines the degree and direction of endogeneity; the case $\rho=0$ corresponds to an exogenous product attribute. Next, given that the variation of $x_{jt}^{unscaled}$ depends on $\rho$, we apply min-max normalization to define $x_{jt}^{endo}=\frac{x_{jt}^{unscaled}-x_{t}^{\min}}{x_{t}^{\max}-x_{t}^{\min}}$, where $x_{t}^{\min}$ and $x_{t}^{\max}$ are the minimum and maximum values of $x_{jt}^{unscaled}$ in market $t$. In this way, we fix the support of the endogenous attribute: $x_{jt}^{endo}\in[0,1]$ for any $\rho$.
@@
Each Monte Carlo configuration is represented by a unique five-tuple $(J,F,T,\phi,F_{c})$. Specifically, $J=36$, $F\in\{1,2,\ldots,36\}$, $T\in\{10,100\}$, $\phi\in\{0,0.1,\ldots,1\}$, and $F_{c}\in\{1,2,\ldots,F-1\}$. [^19] When the DGP includes an endogenous product attribute, $\phi$ is fixed at 1, and $\rho$ is varied to generate unique five-tuples $(J,F,T,\rho,F_{c})$. For each configuration, we generate 500 simulated datasets ($S=500$) to evaluate the finite-sample performance of the proposed test statistic, $T_{IV}^{RV}$, and compare it with the performance of the test based on equilibrium markup conditions, $T_{markup}^{RV}$. We also investigate the performance of $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as instruments for demand estimation. Furthermore, in Section 7, we focus on cases with varying $F$ and fixed $J$ but without collusive behavior ($\phi=0$ and $F_{c}=F$) to evaluate the performance of BLP-type IVs, following the approach of armstrong2016large.
@@ 5.2 Computation of test statistics
For each dataset with configuration $(J,F,T,\phi,F_{c})$, we construct $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as outlined in Section 4.1.202020Note that the summation of the constant term, often referred to as the product counts IV, cannot be utilized in our DGP setting because it fails to generate cross-firm or cross-market variations in the number of products. Moreover, the inclusion of second-order polynomial IVs in (9) ensures that the rank condition for the GMM framework is satisfied in our DGP when a random coefficient is incorporated into the indirect utility function (16). Note that even when $\phi=0$, we still construct $\mathbf{z}^{coll}$ based on a scenario in which the researcher falsely suspects collusion among the firms of interest. To assess the relevance of these two instruments under a given true conduct model, we compute two F-statistics associated with (i) $\bm{\theta}^{1}=0$ (Fstat1) and (ii) $\bm{\theta}^{2}=0$ (Fstat2) from equations (10) and (11), respectively. Our proposed test statistic, $T_{IV}^{RV}$, is then calculated using equation (12) and the description provided therein. Heteroskedasticity-robust standard errors are used to compute $\hat{\sigma}_{IV}$ in equation (12).
@@
as excluded instruments $\mathbf{z}_{jt}=\left(\mathbf{z}_{jt}^{own},\mathbf{z}_{jt}^{other}\right)$. Note that these instruments are constructed based on observed ownership rather than suspected ownership. Next, we specify the marginal cost function as a linear function of the constant term and $x_{jt}$ only, even though it is also linear in a cost shifter $w_{jt}$, as shown in equation (17). [^21] Accordingly, we do not include $w_{jt}$ as an instrument when estimating the demand. Thus, we consider a scenario in which the cost function is potentially misspecified, either due to data limitations on cost shifters or the omission of a relevant cost shifter. This approach enables us to evaluate the performance of the two testing procedures without using additional data. We then follow the procedure outlined in Section 4.2 to construct $T_{markup}^{RV}$ under two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full profit internalization under industry conduct consistent with the effective firm index used in the DGP). [^22] Although the testing procedure built on equilibrium markup conditions allows us to compare conduct models with other values of $\phi$, such as $\phi=0.2$ versus $\phi=0.8$, the primary interest of researchers and regulatory authorities is often to test full profit internalization ($\phi=1$) against competition ($\phi=0$). Therefore, we compute $T_{markup}^{RV}$ under these two conduct models. The standard error $\hat{\sigma}_{markup}$ in equation (15) is adjusted for two-step demand estimation errors and is heteroskedasticity-robust. The computation of $T_{markup}^{RV}$ is done using PyRVtest (pyrvtest).
6 Monte Carlo evidence: testing performance of $T_{IV}^{RV}$ vs $T_{markup}^{RV}$
@@ 6.1 Power of instruments and estimation performance
We begin by assessing the power of our proposed testing procedure outlined in Section 4.1. To do so, we consider a scenario with four firms, two of which collude ($F=4,F_{c}=3$). We generate Monte Carlo datasets, following the design in Section 5, for various values of $\phi\in[0,1]$. The top panel of Table 1 presents the results when a random coefficient is excluded from the indirect utility function (16) in the DGP, while the bottom panel shows the results when it is included.
@@
Table 1: Comparison of IV performance: $\mathbf{z}^{comp}$ vs $\mathbf{z}^{coll}$ ($F=4$, $F_{c}=3$, $T=100$)
@@
Median values ($S=500$) of % of Fstat2 Median
$\phi$ $|\alpha-\hat{\alpha}|$ $|\sigma_{x}-\hat{\sigma}_{x}|$ $\text{Fstat}_{1}$ $|\alpha-\hat{\alpha}|$ $|\sigma_{x}-\hat{\sigma}_{x}|$ $\text{Fstat}_{2}$ $>$ Fstat1 of $s_{o}$
@@
Notes: The table compares the median absolute errors of the estimated price and nonlinear coefficients, as well as the median F-statistics, across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=4,T=100,\phi,F_{c}=3$), obtained using $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$ as instruments individually. The top panel presents the results when a random coefficient is excluded from the indirect utility function (16) in the DGP, while the bottom panel presents the results when it is included.
Our simulation results presented in Table 1 also show that the median absolute error of the estimated price coefficient $\hat{\alpha}$ decreases significantly in response to a rise in $\phi$ when using $\mathbf{z}^{coll}$ as instruments. As a result, $\mathbf{z}^{coll}$ outperforms $\mathbf{z}^{comp}$ in estimating $\alpha$ when firms collude, even with a low degree of profit internalization ($\phi\geq 0.4$ without a random coefficient and $\phi\geq 0.1$ with a random coefficient).
A similar pattern is observed for the median absolute error of the estimated non-linear coefficient $\hat{\sigma}_{x}$. Moreover, for any $\phi$, the median absolute error obtained using $\mathbf{z}^{coll}$ as instruments is smaller than that obtained using $\mathbf{z}^{comp}$. This finding is not contradictory to our earlier observation that, when the degree of profit internalization is low ($\phi\leq 0.3$), using $\mathbf{z}^{coll}$ instead of $\mathbf{z}^{comp}$ leads to lower first-stage F-statistics. This is because identifying $\sigma_{x}$ requires instruments that are correlated not only with price but also with market shares. [^23] As a simple illustration, consider a nested logit specification, a type of random coefficient model. After inverting a demand system (Berry94), the resulting estimation equation using a standard linear instrumental variables approach is as follows: $\log\frac{s_{jt}}{s_{ot}}=\beta x_{jt}-\alpha p_{jt}+\sigma\log s_{j|g,t}+\xi_{jt},$ where $s_{j|g,t}$ denotes the market share of product $j$ within group $g$. Identifying this random coefficient model requires instruments for both the endogenous price and the endogenous market share $s_{j|g,t}$. See berry2014identification for the instrument requirements under general nonparametric settings. We revisit this point in Section 7, where we discuss the role of own- and other-firm instruments.